OpenStudy (anonymous):

Write x2 - 6x + 7 = 0 in the form (x - a)2 = b, where a and b are integers.

OpenStudy (anonymous):

(x - 4)2 = 3 (x - 3)2 = 2 (x - 2)2 = 1 (x - 1)2 = 4

OpenStudy (anonymous):

@ninjasandtigers one more please :)))

OpenStudy (anonymous):

x^2-6x+7=0 x^2-6x+..=-7 x^2-6x+9=-7+9 (x-3)^2=2

OpenStudy (anonymous):

your like a genius thank you i have one more please

OpenStudy (anonymous):

The students of a school went on a picnic by bus. The number of students in each of x buses was the same. The number of rows of seats in each bus was 1 more than the number of buses. The number of students who sat on each row of seats was 2 more than the number of rows of seats in each bus. The expression below shows the total number of students in all the buses: x(x + 1)(x + 3) How can the total number of rows of seats in all the buses combined be calculated? By multiplying x and (x + 1) By multiplying x and (x + 3) By adding (x + 1) and (x + 3) By adding (x + 1) and (x + 2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

im so confused with this one

OpenStudy (anonymous):

idk this one is confusing me

OpenStudy (anonymous):

i'm also doing a math exam right now and i am about to explode

OpenStudy (anonymous):

ya sameee

OpenStudy (anonymous):

i think it is a because it is saying that there is one more row of seats than there is buses

OpenStudy (anonymous):

ok ill try that thanks

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

wait one more please :)

OpenStudy (anonymous):

Out of the 120 students at a summer camp, 72 signed up for canoeing. There were 23 students who signed up for trekking, and 13 of those students also signed up for canoeing. Use a two-way table to organize the information and answer the following question: Approximately what percentage of students signed up for neither canoeing nor trekking?

OpenStudy (anonymous):

10% 12% 38% 32%

OpenStudy (anonymous):

72+23+13=108 120-108=12 12*100=1200 1200/120=10 10%

OpenStudy (anonymous):

thank you so much and good luck with your exam

OpenStudy (anonymous):

thanks you too :)